Notes on the substantive verb in English

David Pierce

June 28, 2007

Contents

1 Introduction 1

2 Grammar 2

3 English verbs 4

4 be 5

1 Introduction

These notes^{1} are ultimately about the verb that in English is called (to) be. I came to write these notes because of a tendency that seems more prevalent in British than in American English. A supplement to the (British) Guardian Weekly in 2005 contained the following sentence:

[Angel’s] boyfriend, Benson, insisted that she had the baby when she belatedly discovered she was pregnant at six months.^{2}

The word had here is to me quite startling; I would say that Benson insisted that Angel have the baby. As it is, the quotation suggests to me that, when Angel discovered she was pregnant, she already had a baby, and Benson was insisting that this was true. But this is not what was intended: Benson was insisting on a future act, not a present state.

In a word, I think the subjunctive form have should have been used, rather than the indicative form had. In a similar situation, a subjunctive form of be is used, for example, in an American novel of 1927:

1I originally composed these notes in 2005; I have done some editing since.

2Guardian Weekly, July 1–7, 2005: ‘Hear Africa 05’ supplement, p. 3: ‘1—Angel Swartbooi, Born: 02.06.05, South Africa’, by David Beresford and Ellen Elmendorp.

Worse things were said of her, and petitions were afloat that she be locked up.^{3}

If the Guardian writers had written this, they might have said petitions were afloat that she was locked up.

Have and had seem to be different forms of the same word, whereas be and was appear to be completely different words. However, we understand be and was to be versions of the one verb (to) be. How did this come about?

What is the history of be, and how did it come to have so many forms? Accounts of be can be found on the web.^{4} Most of my historical information is second hand, and I try to name my sources. This means there are many references to the original Oxford English Dictionary [OED] in particular; I check these against the web version^{5} of the second edition (1989).

2 Grammar

I shall use standard grammatical terminology; but I want first to say something about what the terms mean, as I see them.^{6} A simple sentence involves

1. a speaker,

2. a listener (or listeners), and

3. (usually) a subject (that is, a real subject, which will be referred to by the grammatical subject).

Some sentences have no subject: such sentences are imperative, as will be discussed below in the context of mood; meanwhile, let us consider a sentence with a subject. A certain part of this sentence is called a (grammatical) subject; this names or otherwise distinguishes the real subject. The grammatical subject also indicates the person of the sentence: that is, it tells something about the relation between the subject, the listener, and the speaker. I am treating person as a feature of the whole sentence, not just of parts of the sentence. The person is called:

1. first, if the speaker and subject are identical;

2. second, if the listener and subject are identical;

3Thornton Wilder, The Bridge of San Luis Rey, p. 15 in the Penguin edition.

4http://www.etymonline.com/columns/tobe.htm, 2005.07.11; http://en.wiktionary.org/wiki/be, 2005.07.12.

5http://dictionary.oed.com/

6How I see them is inspired in part by the manual of ancient Greek by Molin and Williamson that was in use at St John’s College in Annapolis when I was a student there in the 1980s. However, I have not looked at that manual in many years.

3. third, otherwise.

Actually, person is not quite so simple, since a sentence also has a number. In English, the number is:

1. singular, if the subject is conceived as one entity;

2. plural, if more than one.

The subject of a plural sentence can be thought of as a class; then the person of a plural sentence is:

1. first, if the speaker is a member of that class;

2. second, if the listener or listeners are included in the class, but the speaker is not;

3. third, otherwise.

A sentence says something about its subject, namely, something that has a time in relation to the uttering of the sentence; this time-relation is indicated by the tense of the sentence, which can be present or past or (sometimes) future (or sometimes variations on these).

Let us say that what the sentence says about its subject (if it has one) is an event. The tense of the sentence concerns the relation between the event and the uttering of the sentence. This utterance is an interaction between the speaker and listener. The sentence also says something about the event in its relation to the speaker alone. What it says is called the mood (or mode) of the sentence; this mood is, in English:

1. indicative, if the event is real for the speaker;

2. subjunctive, if the event is not real (but rather hypothetical, wished for, and so forth), and there is a subject;

3. imperative, if there is no subject.

By this account, imperative sentences have no tense or person. They have a number, which is singular if the listener is one entity, and plural, if more than one. It may be said that an imperative sentence, in the present sense, is in the second person, since a sentence like ‘God bless you’ is a third-person imperative. However, I shall just refer to such a sentence as subjunctive, in contrast with ‘God blesses you,’ an indicative.

Some verb-forms have no person or number, but do have tense: these are the participles (such as going and gone). Other verb-forms do not even have tense: the infinitive (to go) and the gerund (going).

In English, distinctions of person and number can be indicated by (personal) pronouns; first and second persons must be so indicated. Two numbers and three persons would seem to make six kinds of personal pronoun; however, English has lost its second-person singular, and the former

singular plural
first I we
second (thou) (ye,) you
third she/he/it they

Table 1: Personal pronouns

plural now does duty for both numbers. However, since the old singular retains a vestigial presence in the current language, it is listed parenthetically in Table 1 below.

English is said to be one of the descendents of a common Germanic (or ‘Teutonic’) language, which in turn was one of the descendents of Indo-European (IE); but Germanic and Indo-European were not written down, so words supposed to have belonged to them are sometimes written with asterisks to indicate their hypothetical status.

Indo-European and Germanic verbs seem to have exhibited all of the distinctions of person and number. Verbs of current English do not exhibit these distinctions, and it seems foolish to tabulate their forms as if they did.

I am not going to talk about the compound verbs formed with various auxiliaries. Still, three persons, two numbers, two tenses, and two moods would seem to make 24 possibilities; this is not counting the two numbers for an imperative; the two tenses for a participle; the infinitive; and the gerund; that makes 30 possible forms.

3 English verbs

A regular English verb like walk has four forms—five, if you count an archaic form:

1. second-person singular present indicative (archaic): walkest;

2. third-person singular present indicative: walks;

3. present participle, and gerund: walking;

4. past: walked;

5. other: walk.

I am ignoring the archaic alternative third-person singular walketh. I am not treating the infinitive phrase to walk separately from walk.

I do not know why, in the current language, with all of the distinctions possible among the present indicative, subjunctive, and imperative verbs, the third-person singular indicative alone retains a special form. This form can be usefully contrasted with the subjunctive:

1. ‘He insists that he walk’—he will not accept a ride.

2. ‘He insists that he walks’—he denies taking taxis. But the subjunctive is obsolescent. (The first sentence does not sound so natural to me as the equivalent ‘He insists on walking.’) In some verbs, the past participle is distinct from the other past forms, and there are five forms in all: sings, singing, sang, sung, sing. Other verbs collapse distinctions, showing only three forms: puts, putting, put. One verb represents the coalescence of two from the ancestral language, but still exhibits only five forms: goes, going, went, gone, go.

4 be

In a class by itself is be, with eight forms currently, and one more that is archaic:

1. present:

(a) indicative:

i. singular: A. first person: am B. (second person: art) C. third person: is

ii. plural: are

(b) subjunctive (and imperative, and infinitive): be (c) participle (and gerund): being

2. past:

(a) first- and third-person singular indicative: was (b) other indicative, and subjunctive: were

(c) participle: been By the account of [OED] and [HMK], these are the remnants of three verbs in Indo-European:

1. From *(e)s- come am, is, are;

2. from *wes- come was, were;

3. from *bheu- come be, being, been.

present indicative:  eom  am  bēo  be
eart  art  bist beest
is  is  biþ beeth
sind(on), earun are  bōþ  be
present subjunctive  sie  bēo  be
sīen  bēon  be
imperative:  wes  bēo  be
wesaþ  bēoþ  be
infinitive:  wesan  bēon  be
present participle:  wesende  bēonde being
past indicative:  wæs  was
wǣre  wert
wǣron  were
past subjunctive wǣre were
wǣren  were

Table 2: Forms of the substantive verb in Old and modern English

A more recent authority [COD] traces are to a different source; on the web^{7} the source is said to be IE *her-.

By the account of [OED], es- was the ‘Aryan’ (that is, Indo-European) substantive verb, but it had only present forms in ‘Teutonic’ (Germanic) languages like English. In another Germanic language, Gothic, wes- was a separate complete verb with the meaning of ‘remain, stay,’ and hence ‘continue to be.’ Old English (OE, that is, English before the Norman Invasion of 1066) took the past forms of wes- and treated them as past forms of es-, thus creating a complete verb that we might call am-was. Meanwhile, in Old English, be was a distinct verb with no past forms, with the meaning of ‘become, come to be’; hence its present forms could (and did) serve as a future tense for am-was. Later, by the beginning of the 13th century, forms of be supplanted some forms of am-was. Eventually, a past participle from be was introduced (there was none in Old English).

The following Old-English forms are from [HMK]; many variants are given in [OED]. In brackets I give the modern form of the word, if it exists, followed by cognate forms in Greek and Latin (because they are familiar to me), though not in other Germanic tongues (because they are not familiar to me). The forms themselves are summarized in Table 2.

1. Forms from es-:

(a) present indicative:

i. singular:

7http://en.wikipedia.org/wiki/Indo-European_copula, 2005.07.11

A. first: eom [am,^{8} ����� svm]

B. second: eart [art, ��� es]^{9}

C. third: is [is, ����� est]^{10}

ii. plural (all persons):

A. from ’s-: sind(on),^{11} sint [����� svnt]^{12}

B. from es-: earun^{13} [are, ������ ����� �¯���� estis]^{14}

(b) present subjunctive^{15}:

i. singular (all persons^{16}): sie [sīm, sīs, sīt]

ii. plural: sīen [sīmvs, sītis, sīnt]^{17}

2. Forms from wes-:

(a) imperative:

i. singular: wes

ii. plural: wesaþ

(b) infinitive: wesan

(c) present participle: wesende

(d) past indicative:

i. first- and third-person singular: wæs [was^{18}]

ii. second-person singular: wǣre [were^{19}]

8This is the sole survival in English of the personal suffix -m [OED]. 9According to [OED], the s in the root became r, and the t represents the second-person

pronoun; but this formation is not found in Gothic, or outside Germanic. However, art is traced in [COD] to the hypothetical Germanic base *ar-.

10 Since Middle English, is has been used in the north of England for all singular persons, and in the plural, when the subject is a noun or a relative pronoun; such usage is frequent in Shakespeare’s folio of 1613 [OED].

11 The -on is a second plural suffix, occurring in West Germanic [OED]. 12 The Greek ���� is not mentioned in [OED], but according to [Smyth], it represents

����� and can be compared with svnt—which [OED] does mention. Both of these verbs are third person.

13 I do not find this word in [HMK]; says [OED], it was a re-formation from the full es-, but was used only in Anglian dialects, alongside the formation on ’s-.

14 The persons of the Greek and Latin forms are, respectively, first, second, Homeric third [Smyth], and second.

15 The subjunctive was originally an optative [OED]; see Note 17. 16 A form peculiar to the first person was possible: siēm or siōn [OED]. 17 To sīmvs can be compared the optative ������� from ���¯���� [Smyth]. 18 In dialect, were and war are used [OED]. 19 This is the etymological form; the analogical form wast is found in Tindale and the

1611 Bible; in Shakespeare is the intermediate form wert [OED].

iii. plural: wæron [were^{20}]

(e) past subjunctive:

i. first- and third-person singular: wǣre [were]

ii. second-person singular: wǣre [wert^{21}]

iii. plural: wǣren [were^{22}]

3. Forms from be:

(a) present indicative:

i. singular^{23}:

A. first person: bēo^{24} [be, ���� fui]

B. second person: bist^{25} [beest]

C. third person: biþ^{26} [beeth, bes]

ii. plural: bēþ [be^{27}]

(b) present subjunctive^{28}:

i. singular: bēo [be^{29}]

ii. plural: bōn [be]

(c) imperative:

i. singular: bēo [be]

ii. plural: bēoþ [be]

20 In the 16th–18th cc., was was used almost universally with you, if this was used for one listener [OED].

21 Formerly were; the -t is by analogy with the indicative; also, was was common in the 17th–18th cc., when this was even used as a plural form by those who used it also as an indicative plural [OED].

22 See Note 21. 23 These forms are said to be regular in the dialect of southern England [OED]. 24 Usually with future meaning [OED]. 25 Usually with future meaning [OED]. 26 Future or present meaning [OED]. 27 Before 1250, this replaced sind(on) in the South and became standard for centuries;

we see it in ‘the powers that be’ [OED].

28 Indicative forms have been substituted for imperative occasionally since the 15th c., and are now usually used after if, though, unless, etc.; consider the 1611 Bible, I John 4:1: ‘Try the spirits whether they are of God’ [OED].

29 In the second person, after if, though, etc., beest was common in thte 16th–17th cc. [OED]. Indeed, consider John Donne:

If thou beest born to strange sights, Things invisible to see, Ride ten thousand days and nights, Till age snow white hairs on thee…

and note that ‘snow’ here is a subjunctive verb.

(d) infinitive: i. bēon [be] ii. to bēonne [to be]

(e) present participle: bēonde [being]

(f) past participle: (ben(e)^{30}) [been^{31}]

References

[COD] T. F. Hoad, editor. The Concise Oxford Dictionary of English Etymology. Oxford University Press, Oxford, 1986.

[OED] Murray et al., editors. The Compact Edition of the Oxford English Dictionary. Oxford University Press, 1973.

[HMK] James R. Hulbert Samuel Moore, Thomas A. Knott. The Elements of Old English: Elementary Grammar, Reference Grammar, and Reading Selections. George Wahr Publishing Co., Ann Arbor, Michigan, tenth edition, 1977. Revised, Enlarged and Corrected.

[Smyth] Herbert Weir Smyth. Greek Grammar. Harvard University Press, Cambridge, Massachussets, 1980. Revised by Gordon M. Messing. Eleventh Printing.

30 Not known in Œ.

31 In 14th–15th cc., be was common in literature [OED].